OpenStudy (anonymous):

how do I find the complex roots of the function x^4 + 7x^2 + 6 = 0

OpenStudy (anonymous):

use the quadratic formula to solve for \(x^2\) then take the square root of your answers

OpenStudy (anonymous):

actually this one factors as \[(x^2+6)(x^2+1)=0\]

OpenStudy (anonymous):

so just forget about the x^4 and plug the rest in to the quadratic equation?

OpenStudy (anonymous):

telling you \[x^2=-6\implies x=\pm\sqrt{6}i\] or \[x^2=-1\implies x=\pm i\]

OpenStudy (anonymous):

if it did not factor so easily, you would use the quadratic formula, yes

OpenStudy (anonymous):

You factored it just by intuition or what?